Assume that GDP (Y) IS 5,000. Consumption (C) is given by the equation C-1,200+0.5(Y-T)-50r, where r is the real interest rate in percentage. Investment (I) is given by the equation I=1,500-50r. Taxes (T) are 1,200 and government spending (G) is 1,500.

1) What are the equilibrium values of C, I, and r?

2) What are the values of private saving, public saving, and national saving?
Answer #1

In Equilibrium Y = c+Ith y = 1200 +08 (Y-T) - Sopot isoo - SOP + 1 Soo y = 4200 tos [y - 1200] -100 P y = 4200 toisy - 600 -C = 12oo tois (sooo -1200) - SSO C = 1200 + 1900-550 = $25so Substitating. I value in I = 1500-S0 (1) I equation =1500-SSO (3